# Homogenization framework for rigid and non-rigid foldable origami metamaterials

### v1

@@ Abstract
Origami metamaterials typically consist of folded sheets with periodic patterns, conferring them with remarkable mechanical properties. In the context of Continuum Mechanics, the majority of existing predictive methods are mechanism analogs which favor rigid folding and panel bending. While effective in predicting primary deformation modes, existing methods fall short in capturing the full spectrum of deformation of non-rigid foldable origami, such as the emergence of curvature along straight creases, local strain at vertices and warpage in panels. To fully capture the entire deformation spectrum and enhance the accuracy of existing methods, this paper introduces a homogenization framework for origami metamaterials where the faces are modeled as plate elements. Both asymptotic and energy-based homogenization methods are formulated and implemented. As a representative crease pattern, we examine the Miura origami sheet homogenized as an equivalent Kirchhoff-Love plate. The results reveal that certain effective elastic properties are nonlinearly related to both the initial fold angle and the crease stiffness. When benchmarked with results from fully resolved simulations, our framework yields errors up to 12.9%, while existing models, including the bar-and-hinge model and the rigid-panel model, show up to 161% error. The differences in errors are associated with the complex modes of crease and panel deformation in non-rigid origami, unexplored by the existing models. This work demonstrates a precise and efficient continuum framework for origami metamaterials as an effective strategy for predicting their elastic properties, understanding their mechanics, and designing their functionalities.
@@ 1 Introduction
Origami, the art of paper folding, enables the folding of a flat, thin sheet into a three-dimensional structure with remarkable kinematics and mechanical properties. Initially considered as an art, origami has so far inspired the development of multiple technologies for applications, such as soft robotic actuators (Chi et al., 2022; Li et al., 2019b; Yang et al., 2024; Xue et al., 2024), deployable structures (Dang et al., 2025; Jamalimehr et al., 2022; Overvelde et al., 2017; Melancon et al., 2021; Fang et al., 2024; Li et al., 2023), energy absorbers (Jiang et al., 2025; Almessabi et al., 2024), and foldable electronics (Zhang et al., 2022; Li et al., 2019a). Crease patterns with period tessellation enable the design of mechanical metamaterials, namely origami metamaterials, whose physical properties are primarily governed by the architecture of the pattern rather than the base material. Figure 1 shows some examples of origami metamaterials. A wide range of properties have been investigated thus far, such as energy absorption capacity (Gattas and You, 2015; Harris and McShane, 2021), wave propagation (Oudghiri-Idrissi and Guzina, 2022; Zhang and Rudykh, 2024; Yasuda et al., 2019), and tunable or re-programmable geometry (Mirzajanzadeh and Pasini, 2025; Nassar et al., 2022; Castle et al., 2016; Gao et al., 2023).
@@
Despite the wealth of research in origami metamaterials, modeling their mechanical behavior still poses significant challenges. One of the earliest approaches, the rigid-panel model (Schenk and Guest, 2013; Wei et al., 2013; Hu and Liang, 2020), treats an origami metamaterial as an assembly of rigid panels connected with revolute joints forming a mechanism. The kinematics can thus be studied through the theory of mechanisms, and the structural mechanics through the force-displacement relation. Later, the bar-and-hinge model (Schenk and Guest, 2011; Filipov et al., 2017; Liu and Paulino, 2017; Overvelde et al., 2017) was proposed to include the role of panel deformation. The origami metamaterial was thus treated as a combination of a truss lattice and rotational springs (Lahiri and Pratapa, 2023). The former allows both extension in truss elements and revolution around pin joints, capturing the primary deformation modes at the global level. For non-rigid foldable origami patterns, another popular option is to model their panels with plate Finite Elements (FE) (Hu et al., 2021; Filipov et al., 2015; Feng et al., 2022). This method gives an accurate description of the stress and strain distribution but is computationally demanding. The computational requirement becomes overwhelming when there is a large origami tessellation with a complex fold pattern. For this reason, the FE analysis of origami specimens is limited to a few unit cells of relatively simple patterns.
@@
To bypass the fully resolved, detailed simulation of a periodic metamaterial, homogenization is often used to approximate the global response using an equivalent continuum. Multiple theories exist for homogenization, and are effective in simplifying the mechanical analysis of large, periodic tessellations, such as cellular origami metamaterials (Cheung et al., 2014), origami sheets (Heimbs, 2013) and origami tubes (Turco et al., 2024), corresponding to 3D, 2D and 1D tessellations respectively, for both the in-plane and out-of-plane mechanical properties. Cheung et al. (2014) applied periodic boundary conditions to an interleaved tube cellular structure to examine its elastic properties with classical dimensional scaling analysis. Vasudevan and Pratapa (2024) presented an energy-based homogenization framework to determine the effective elastic constants of a couple-stress continuum representing origami sheets. Nassar et al. (2017) leveraged asymptotic homogenization to fit smooth surfaces with origami tessellations. Regardless of the method, the idea of homogenization proves to be a powerful tool in revealing the unique kinematics and mechanics of origami metamaterials.
@@
Existing methods to homogenize the mechanical properties of origami metamaterials mainly capture rigid-foldable modes and rely on specific assumptions for folding. One is that creases and panels have contrasting properties, enabling the metamaterial to fold as a mechanism. Another is to treat the metamaterial as a non-reconfigurable structure that deforms in the panels rather than folding, similarly to a foldcore sandwich plate designed for load bearing and energy absorption (Zheng et al., 2022; Sturm et al., 2015; Zhang et al., 2021). Most works adopt either of these assumptions as they are easily implemented with the modeling methods reviewed above. However, origami systems may encompass a wide range of panel and crease characteristics. An example is soft origami actuators made of elastomer (Martinez et al., 2012; Jin et al., 2022). The soft and stretchable base material allows panels to bulge and wrinkle. Describing such mechanical responses is challenging for the rigid panel model and the bar-and-hinge model. At the same time, elastomeric creases may not show contrasting mechanical properties compared to the surrounding panels. Although this does not prevent the origami structure from folding into a wide range of motion, the assumption of rigid folding no longer applies. This is one among several other examples that call for the development of a property prediction method which is compatible with a broad variety of attributes in panels and creases so as to capture the mechanics of non-rigid foldable origami metamaterials.
@@ 2 Homogenization formulations for origami metamaterials
With a focus on a single origami sheet, we present the homogenization formulations for determining the effective elastic constants of an equivalent Kirchhoff-Love plate. A common approach of all homogenization methods is to take a representative region, which in our paper is a unit cell, and assume that its properties are equivalent to those of the entire metamaterial (Hollister and Kikuchi, 1992). This section presents two formulations, i.e., asymptotic and energy-based methods. They tackle the homogenization problem from dissimilar perspectives. The former approximates the field variables in the governing equation, such as the displacement, by an asymptotic expansion. The latter, on the other hand, approximates the in-situ stress or strain of the metamaterial with applied boundary conditions of a unit cell. By comparing both methods, we can appreciate their common and distinct traits, obtain the effective properties, and identify the strain energy density of the metamaterial under prescribed deformation modes.
@@ 2.1 Homogenized Kirchhoff-Love plate model for origami sheets
For homogeneous or laminated plates, there exist relationships between the stiffness coefficients $A^{H}_{\alpha\beta\gamma\delta}$, $B^{H}_{\alpha\beta\gamma\delta}$, $D^{H}_{\alpha\beta\gamma\delta}$ and the base material properties obtained from the Cauchy continuum theory (Reddy, 2006). However, this simplification does not apply to a nonhomogeneous medium such as a origami metamaterial with a general fold pattern. Rather, $A^{H}_{\alpha\beta\gamma\delta}$, $B^{H}_{\alpha\beta\gamma\delta}$, and $D^{H}_{\alpha\beta\gamma\delta}$ should all be computed independently (Vasudevan and Pratapa, 2024). The strain energy density of a homogeneous thin plate is
@@ 2.2.1 Theory
Consider the representative origami sheet shown in Figures 3 and 3. $\Omega_{h}$ denotes the domain of the sheet. The upper, lower, and lateral surfaces are respectively $\Gamma_{h}^{+},\Gamma_{h}^{-},\Gamma_{h}^{\textup{lat}}$. A small parameter $h$ denotes the domain height of the equivalent homogenized plate. There are $n$ unit cells along one given side of the sheet, with a total length of $\ell$. The origami pattern is periodic with a unit cell size of $\varepsilon=\ell/n$. Since the origami sheet is single-layered, its height $h$ and unit cell size $\varepsilon$ have the same order of magnitude. They are small parameters, meaning $h\to 0$ and $\varepsilon\to 0$ simultaneously as the origami tessellation becomes infinitely large.
@@
where the superscripts $h,\varepsilon$ signifies that the elastic stiffness is a function of the height and the periodicity. The indices $i,j,k,\ell$ range from 1 to 3. The following change of variables is now adopted to simplify the analysis: all functions of the origami sheet are made independent of $h$, i.e., $z_{1}=x_{1}$, $z_{2}=x_{2}$, and $z_{3}=x_{3}/h$. This results in a domain $\Omega$ for the origami sheet, an upper surface $\Gamma^{+}$, a lower surface $\Gamma^{-}$, and a lateral surface $\Gamma^{\textup{lat}}$. The elastic stiffness $C_{ijk\ell}^{h\varepsilon}(x_{1},x_{2},x_{3})$ becomes $C_{ijk\ell}^{\varepsilon}(z_{1},z_{2},z_{3})$, which is a function of the periodicity $\varepsilon$ but not a function of the height $h$. The full problem of elasticity of the origami sheet is
@@
where each term $u_{i}^{(0)}$, $u_{i}^{(1)}$… are functions of both coordinates $\mathbf{x}$ and $\mathbf{y}$. These terms are smooth in $\mathbf{x}$ and periodic in $y_{1}$ and $y_{2}$. On the other hand, the stress tensor expands into
@@
where $L_{1}$ and $L_{2}$ are the lengths of the unit cell in the $y_{1}$ and $y_{2}$ directions, respectively. Equation 11 applies to unit cells with a general shape. Since an origami unit cell consists of thin panels, we simplify the expression by assuming that the panels are Kirchhoff-Love plates. Their collective mid-surface $Y^{0}$ is shown in Figure 4. $Y^{0}$ is referred to the coordinates $(y_{1}^{\prime},y_{2}^{\prime})$. Note that the plates in a unit cell are not the homogenized plate introduced in Section 2.1. With a displacement field $\mathbf{w}$, the local strain field in the unit cell panels is $\mathbf{e}=\boldsymbol{\mu}+y_{3}\boldsymbol{\kappa}$ with
@@
where $\mathbf{e}^{0\kappa,(\alpha\beta)}$ denotes unit curvature or twist in direction $\alpha\beta$, $\boldsymbol{\mu}^{*\kappa,(\alpha\beta)}$ is the strain of characteristic displacement $\mathbf{p}^{\alpha\beta}$ from the second microscopic problem in Equation 10, and $\boldsymbol{\kappa}^{*\kappa,(\alpha\beta)}$ is the curvature of the characteristic displacement $\mathbf{p}^{\alpha\beta}$. The unit cell strain energy $U$ is calculated from the sum of unit curvature $\mathbf{e}^{0\kappa,(\alpha\beta)}$ and the strain of corresponding characteristic displacement $\boldsymbol{\mu}^{*\kappa,(\alpha\beta)}+y_{3}\boldsymbol{\kappa}^{*\kappa,(\alpha\beta)}$. This strain energy relates to the diagonal term as
@@
where $\Omega^{0}$ is the mid-surface of the homogenized plate, $\Gamma^{0}$ is the boundary of the mid-surface, $N_{\alpha\beta}$ are the force resultants, $M_{\alpha\beta}$ are the moment resultants, $T_{1}$, $T_{2}$, and $T_{3}$ are averaged tractions defined as $T_{i}=\frac{h}{|Y|}\int_{\Gamma^{\pm}}t_{i}^{\pm}\,ds$, $Q_{\alpha}$ are averaged bending moments defined as $Q_{\alpha}=\frac{h}{|Y|}\int_{\Gamma^{\pm}}y_{3}t_{\alpha}^{\pm}\,ds$, $\bar{u}_{\alpha}$ are the homogenized displacements, and $u^{(0)}_{3}$ is the homogenized deflection.
@@ 2.2.2 Numerical implementation
Given the geometry of origami patterns, which can be complex, it is often not feasible to use closed-form expressions to perform asymptotic homogenization. Rather, a numerical approach becomes preferable. Our method here relies on FE analysis to find numerical solutions to the microscopic problems (Equations 9 and 10) and the effective stiffness coefficients (Equations 14, 19 and 23). The following procedure is fully implemented through ABAQUS python scripting, with python codes and input files available at https://github.com/Wendyli119/OMhomogenization.git. However, it is possible to generalize the procedure to any FE software package with minor modifications. Our implementation differs from the literature (Cai et al., 2014; Eskandari et al., 2024) in the strain energy calculation. Rather than calculating the strain energy from the nodal quantities during postprocessing, this step is performed fully using the selected FE software. This modification adds flexibility to the choice of element types. It is relevant to models of origami tessellations that require various types of plate elements, shell elements and torsional springs.
@@
Panel side length $a$ ($m$) 20
Panel side length $b$ ($m$) 20
Sector angle $\gamma$ () 60
Initial fold angle $\theta_{0}$ () 30
Panel thickness $t$ ($m$) 0.13
Unit length crease stiffness $K_{\textup{cr}}/a$ (N) 0.1
Base material Young’s modulus $E$ (MPa) 4000
@@
Figure 6: Step 1 of asymptotic homogenization showing six of the deformation modes with unit strain or unit curvature. To ensure uniform strain or curvature distribution, displacements are prescribed nodally as shown in the inset. Displacements are scaled.
@@
where $\mathbf{f}^{0}$ is the initial strain loading obtained from the previous step, and $\boldsymbol{\chi}^{*}$ is the unknown characteristic displacement field. Different from the unit strains and curvatures, the characteristic displacements are assumed to be periodic in $y_{1}$ and $y_{2}$, as specified in Equations 9 and 10. This means that opposite unit cell boundaries have identical displacement. Because the Miura pattern tessellates in the $(y_{1},y_{2})$ directions, each unit cell has four boundaries, with the right boundary connected to the left boundary of an adjacent unit cell, and similar for the top and bottom boundaries. Accordingly the periodic boundary condition constrains the displacements of each pair of boundary nodes as
@@
Figure 7: Step 2 of asymptotic homogenization solving the microscopic problem. (a) Periodic boundary conditions are applied on the boundary displacements according to Equation 30. (b) Numerical solutions of the characteristic displacement field $\boldsymbol{\chi}^{*}$. Displacements are scaled.
@@
Figure 8: Step 3 of asymptotic homogenization computing unit cell strain energy $U$. (a) Difference between the unit strain displacement field $\boldsymbol{\chi}^{0}$ and the characteristic displacement field $\boldsymbol{\chi}^{*}$. Only 6 of the deformation modes are shown with lighting effects in 3D rendering. Displacements are scaled. (b) Unit cell strain energy $U$. Colors represent the strain energy magnitudes of the corresponding deformation modes.
@@
The final step is the homogenized problem. The origami sheet is simplified into a thin plate. The homogenized plate $\Omega$, shown in Figure 2, is modeled with triangular or quadrilateral thin plate elements with a plate thickness of $h$. The effective elastic properties are specified in the FE model as elastic and anisotropic material properties. After applying appropriate loads and boundary conditions on the homogenized model, one last static analysis is performed, which solves the homogenized governing equation Equation 25.

### v3

@@ Abstract
[…]
@@ 1 Introduction
Origami, the art of paper folding, enables the folding of a flat, thin sheet into a three-dimensional structure with remarkable kinematics and mechanical properties. Initially considered as an art, origami has so far inspired the development of multiple technologies for applications, such as soft robotic actuators (Chi et al., 2022; Li et al., 2019b; Yang et al., 2024; Xue et al., 2024), deployable structures (Dang et al., 2025; Jamalimehr et al., 2022; Overvelde et al., 2017; Melancon et al., 2021; Fang et al., 2024; Li et al., 2023), energy absorbers (Jamalimehr et al., 2026; Jiang et al., 2025; Almessabi et al., 2024), and foldable electronics (Zhang et al., 2022; Li et al., 2019a). Crease patterns with period tessellation enable the design of mechanical metamaterials, namely origami metamaterials, whose physical properties are primarily governed by the architecture of the pattern rather than the base material. Figure 1 shows some examples of origami metamaterials. A wide range of properties have been investigated thus far, such as energy absorption capacity (Gattas and You, 2015; Harris and McShane, 2021), wave propagation (Oudghiri-Idrissi and Guzina, 2022; Zhang and Rudykh, 2024; Yasuda et al., 2019), and tunable or re-programmable geometry (Mirzajanzadeh and Pasini, 2025; Nassar et al., 2022; Castle et al., 2016; Gao et al., 2023).
@@
Despite the wealth of research in origami metamaterials, modeling their mechanical behavior still poses significant challenges. One of the earliest approaches, the rigid-panel model (Schenk and Guest, 2013; Wei et al., 2013; Hu and Liang, 2020), treats an origami metamaterial as an assembly of rigid panels connected with revolute joints forming a mechanism. The kinematics can thus be studied through the theory of mechanisms, and the structural mechanics through the force-displacement relation. Later, the bar-and-hinge model (Schenk and Guest, 2011; Filipov et al., 2017; Liu and Paulino, 2017; Overvelde et al., 2017) was proposed to include the role of panel deformation. The origami metamaterial was thus treated as a combination of a truss lattice and rotational springs (Lahiri and Pratapa, 2023). The former allows both extension in truss elements and revolution around pin joints, capturing the primary deformation modes at the global level. The bar-and-hinge model has been used to capture the overall shape change under large deformation (Xu et al., 2024, 2025). However, local stress and strain are often neglected or simplified by this method. For non-rigid foldable origami patterns, another popular option is to model their panels with plate Finite Elements (FE) (Hu et al., 2021; Filipov et al., 2015; Feng et al., 2022). This method gives an accurate description of the stress and strain distribution but is computationally demanding. The computational requirement becomes overwhelming when there is a large origami tessellation with a complex fold pattern. For this reason, the FE analysis of origami specimens is limited to a few unit cells of relatively simple patterns.
@@
To bypass the fully resolved, detailed simulation of a periodic metamaterial, homogenization is often used to approximate the global response using an equivalent continuum. Multiple theories exist for homogenization, and are capable of simplifying the mechanical analysis of large, periodic tessellations, such as cellular origami metamaterials (Cheung et al., 2014; Lyu et al., 2021), origami sheets (Heimbs, 2013) and origami tubes (Turco et al., 2024), corresponding to 3D, 2D and 1D tessellations, respectively, for both the in-plane and out-of-plane mechanical properties. Xu et al. (2024, 2025) proposed an effective surface theory and its numerical implementation on nonlinear homogenization to characterize origami soft modes using the bar-and-hinge model. Cheung et al. (2014) applied periodic boundary conditions to an interleaved tube cellular structure to examine its elastic properties with classical dimensional scaling analysis. Vasudevan and Pratapa (2024) presented an energy-based homogenization framework with bar-and-hinge to determine the effective elastic constants of a couple-stress continuum representing origami sheets. Nassar et al. (2017) leveraged asymptotic homogenization to fit smooth surfaces with bar-and-hinge origami tessellations. Regardless of the method, the idea of homogenization proves to be a powerful tool in revealing the unique kinematics and mechanics of origami metamaterials.
@@
Existing methods to homogenize the mechanical properties of origami metamaterials mainly capture rigid-foldable modes and rely on specific assumptions for folding. One is that creases and panels have contrasting properties, enabling the metamaterial to fold as a mechanism. Another is to treat the metamaterial as a non-reconfigurable structure that deforms in the panels rather than folding, similarly to a foldcore sandwich plate designed for load bearing and energy absorption (Zheng et al., 2022; Sturm et al., 2015; Zhang et al., 2021). Most works adopt either of these assumptions as they are easily implemented with the modeling methods reviewed above. However, origami systems may encompass a wide range of panel and crease characteristics. An example is soft origami actuators made of elastomer (Martinez et al., 2012; Jin et al., 2022). The soft and stretchable base material allows panels to bulge and wrinkle. Describing such mechanical responses is challenging for the rigid panel model and the bar-and-hinge model. At the same time, elastomeric creases may not show contrasting mechanical properties compared to the surrounding panels. Although this does not prevent the origami structure from folding into a wide range of motion, the assumption of rigid folding no longer applies. This is one among several other examples that call for the development of a property prediction method which is compatible with a broad variety of attributes in panels and creases so as to capture the mechanics of non-rigid foldable origami metamaterials. To address this gap, we investigate the previously uncharacterized deformation modes of origami metamaterials under varying crease stiffness.
@@ 2 Homogenization formulations for origami metamaterials
With a focus on a single origami sheet, we present the homogenization formulations for determining the effective elastic constants of an equivalent Kirchhoff-Love plate. A common approach of all homogenization methods is to take a representative region, which in our paper is a unit cell, and assume that its properties are equivalent to those of the entire metamaterial (Hollister and Kikuchi, 1992). This section presents two formulations, i.e., asymptotic and energy-based methods. They tackle the homogenization problem from dissimilar perspectives. The former approximates the field variables in the governing equation, such as the displacement, by an asymptotic expansion. The latter, on the other hand, approximates the in-situ stress or strain of the metamaterial with applied boundary conditions of a unit cell. Both methods are applicable to origami metamaterials under linear elastic deformation. The pattern should exhibit translational periodicity, with the unit cell size much smaller than the full tessellation. Finally, the unit cell height, length, and width should be on the same order of magnitude. By comparing both methods, we can appreciate their common and distinct traits, obtain the effective properties, and identify the strain energy density of the metamaterial under prescribed deformation modes.
@@ 2.1 Homogenized Kirchhoff-Love plate model for origami sheets
For homogeneous or laminated plates, there exist relationships between the stiffness coefficients $A^{H}_{\alpha\beta\gamma\delta}$, $B^{H}_{\alpha\beta\gamma\delta}$, $D^{H}_{\alpha\beta\gamma\delta}$ and the base material properties obtained from the Cauchy continuum theory (Reddy, 2006). However, this simplification does not apply to a nonhomogeneous medium such as an origami metamaterial with a general fold pattern. Rather, $A^{H}_{\alpha\beta\gamma\delta}$, $B^{H}_{\alpha\beta\gamma\delta}$, and $D^{H}_{\alpha\beta\gamma\delta}$ should all be computed independently (Vasudevan and Pratapa, 2024). The strain energy density of a homogeneous thin plate is
@@ 2.2.1 Theory
Consider the representative origami sheet shown in Figures 3 and 3. $\Omega_{h}$ denotes the domain of the sheet. A small parameter $h$ denotes the domain height of the equivalent homogenized plate. The upper, lower, and lateral surfaces are respectively $\Gamma_{h}^{+}$, $\Gamma_{h}^{-}$, $\Gamma_{h}^{\textup{lat}}$. $\Gamma_{h}^{+}$ and $\Gamma_{h}^{-}$ consist of all points $(x_{1},x_{2},x_{3})$ such that $(x_{1},x_{2})$ lies in the mid-surface $\Omega^{0}$. On $\Gamma_{h}^{+}$ $x_{3}=h$, while on $\Gamma_{h}^{-}$ $x_{3}=0$. $\Gamma_{h}^{\textup{lat}}$ has $(x_{1},x_{2})$ lying on the boundary of $\Omega^{0}$ and $x_{3}$ varying from 0 to $h$. There are $n$ unit cells along one given side of the sheet, with a total length of $\ell$. The origami pattern is periodic with a unit cell size of $\varepsilon=\ell/n$. Since the origami sheet is single-layered, its height $h$ and unit cell size $\varepsilon$ have the same order of magnitude. They are small parameters, meaning $h\to 0$ and $\varepsilon\to 0$ simultaneously as the origami tessellation becomes infinitely large.
@@
where the superscripts $h,\varepsilon$ signify that the elastic stiffness is a function of the height and the periodicity. The indices $i,j,k,\ell$ range from 1 to 3. The following change of variables is now adopted to simplify the analysis: all functions of the origami sheet are made independent of $h$, i.e., $z_{1}=x_{1}$, $z_{2}=x_{2}$, and $z_{3}=x_{3}/h$. This results in a domain $\Omega$ for the origami sheet, an upper surface $\Gamma^{+}$, a lower surface $\Gamma^{-}$, and a lateral surface $\Gamma^{\textup{lat}}$. The elastic stiffness $C_{ijk\ell}^{h\varepsilon}(x_{1},x_{2},x_{3})$ becomes $C_{ijk\ell}^{\varepsilon}(z_{1},z_{2},z_{3})$, which is a function of the periodicity $\varepsilon$ but not a function of the height $h$. The full problem of elasticity of the origami sheet is
@@
where each term $u_{i}^{(0)}$, $u_{i}^{(1)}$… are functions of both coordinates $\mathbf{x}$ and $\mathbf{y}$. These terms are smooth in $\mathbf{x}$ and periodic in $y_{1}$ and $y_{2}$. On the other hand, the stress tensor expands to
@@
where $L_{1}$ and $L_{2}$ are the lengths of the unit cell along $y_{1}$ and $y_{2}$, respectively. Equation 11 applies to unit cells with a general shape. Since an origami unit cell consists of thin panels, we simplify the expression by assuming that the panels are Kirchhoff-Love plates. Their collective mid-surface $Y^{0}$ is shown in Figure 4. $Y^{0}$ is referred to the coordinates $(y_{1}^{\prime},y_{2}^{\prime})$. Note that the plates in a unit cell are not the homogenized plate introduced in Section 2.1. With a displacement field $\mathbf{w}$, the local strain field in the unit cell panels is $\mathbf{e}=\boldsymbol{\mu}+y_{3}\boldsymbol{\kappa}$ with
@@
where $\mathbf{e}^{0\kappa,(\alpha\beta)}$ denotes unit curvature or twist in direction $\alpha\beta$, $\boldsymbol{\mu}^{*\kappa,(\alpha\beta)}$ is the strain of characteristic displacement $\mathbf{p}^{\alpha\beta}$ from the second microscopic problem in Equation 10, and $\boldsymbol{\kappa}^{*\kappa,(\alpha\beta)}$ is the curvature of the characteristic displacement $\mathbf{p}^{\alpha\beta}$. The unit cell strain energy $U$ is calculated from the sum of unit curvature $\mathbf{e}^{0\kappa,(\alpha\beta)}$ and the strain of corresponding characteristic displacement $\boldsymbol{\mu}^{*\kappa,(\alpha\beta)}+y_{3}\boldsymbol{\kappa}^{*\kappa,(\alpha\beta)}$. This strain energy relates to the diagonal terms as
@@
where $\Omega^{0}$ is the mid-surface of the homogenized plate, $\Gamma^{0}$ is the boundary of the mid-surface, $N_{\alpha\beta}$ are the force resultants, $M_{\alpha\beta}$ are the moment resultants, $T_{1}$, $T_{2}$, and $T_{3}$ are averaged tractions defined as $T_{i}=\frac{h}{|Y|}\int_{\Gamma^{\pm}}t_{i}^{\pm}\mathop{\mathrm{d}\!}s$, $Q_{\alpha}$ are averaged bending moments defined as $Q_{\alpha}=\frac{h}{|Y|}\int_{\Gamma^{\pm}}y_{3}t_{\alpha}^{\pm}\mathop{\mathrm{d}\!}s$, $\bar{u}_{\alpha}$ are the homogenized displacements, and $u^{(0)}_{3}$ is the homogenized deflection.
@@ 2.2.2 Numerical implementation
Given the geometry of origami patterns, which can be complex, it is often not feasible to use closed-form expressions to perform asymptotic homogenization. Rather, a numerical approach becomes preferable. Our method here relies on FE analysis to find numerical solutions to the microscopic problems (Equations 9 and 10) and the effective stiffness coefficients (Equations 14, 19 and 23). The following procedure is fully implemented through ABAQUS python scripting, with python codes and input files available (Li, 2025). However, it is possible to generalize the procedure to any FE software package with minor modifications. Our implementation differs from the literature (Cai et al., 2014; Eskandari et al., 2024) in the strain energy calculation. Rather than calculating the strain energy from the nodal quantities during postprocessing, this step is performed fully using the selected FE software. This modification adds flexibility to the choice of element types. It is relevant to models of origami tessellations that require various types of plate elements, shell elements and torsional springs.
@@
Panel side length $a$ $m$ 20
Panel side length $b$ $m$ 20
Sector angle $\gamma$ 60
Initial fold angle $\theta_{0}$ 30
Panel thickness $t$ $m$ 0.13
Unit length crease stiffness $K_{\textup{cr}}/a$ N 0.1
Base material Young’s modulus $E$ MPa 4000
@@
Figure 6: Step 1 of asymptotic homogenization showing six of the deformation modes with unit strain or unit curvature. (a) Displacements are prescribed nodally to ensure uniform strain or curvature distribution. (b) Displacement fields $\boldsymbol{\chi}^{0}$ corresponding to unit strain ($e_{11}=1$ or $e_{22}=1$ or $e_{12}=1$) or unit curvature ($\kappa_{11}=1$ or $\kappa_{22}=1$ or $\kappa_{12}=1$). Displacements are down-scaled for plotting purposes.
@@
where $\mathbf{f}^{0}$ is the initial strain loading obtained from the previous step, and $\boldsymbol{\chi}^{*}$ is the unknown characteristic displacement field. Different from the unit strains and curvatures, the characteristic displacements are assumed to be periodic in $y_{1}$ and $y_{2}$, as specified in Equations 9 and 10. This means that opposite unit cell boundaries have identical displacement. Because the Miura pattern tessellates in $(y_{1},y_{2})$, each unit cell has four boundaries, with the right boundary connected to the left boundary of an adjacent unit cell, and similar for the top and bottom boundaries. Accordingly the periodic boundary condition constrains the displacements of each pair of boundary nodes as
@@
Figure 7: Step 2 of asymptotic homogenization solving the microscopic problem. (a) Periodic boundary conditions are applied on the boundary displacements according to Equation 30. (b) Numerical solutions of the characteristic displacement field $\boldsymbol{\chi}^{*}$. Displacements are down-scaled for plotting purposes.
@@
Figure 8: Step 3 of asymptotic homogenization computing unit cell strain energy $U$. (a) Difference between the unit strain displacement field $\boldsymbol{\chi}^{0}$ and the characteristic displacement field $\boldsymbol{\chi}^{*}$. Only 6 of the deformation modes are shown with lighting effects in 3D rendering. Displacements are down-scaled for plotting purposes. (b) Unit cell strain energy $U$. Colors represent the strain energy magnitudes of the corresponding deformation modes.
@@
The final step is the homogenized problem. The origami sheet is simplified into a thin plate. The homogenized plate $\Omega$, shown in Figure 2, is modeled with triangular or quadrilateral thin plate elements with a plate thickness of $h$. The effective elastic properties are specified in the FE model as elastic and anisotropic material properties. After applying appropriate loads and boundary conditions on the homogenized model, one last static analysis is performed, which solves the homogenized governing equation, see Equation 25.